# InstructAgent: Building User Controllable Recommender via LLM Agent

### v1

@@
Abstract.
Traditional recommender systems usually take the user-platform paradigm, where users are directly exposed under the control of the platform’s recommendation algorithms. However, the defect of recommendation algorithms may put users in very vulnerable positions under this paradigm. First, many sophisticated models are often designed with commercial objectives in mind, focusing on the platform’s benefits, which may hinder their ability to protect and capture users’ true interests. Second, these models are typically optimized using data from all users, which may overlook individual user’s preferences. Due to these shortcomings, users may experience several disadvantages under the traditional user-platform direct exposure paradigm, such as lack of control over the recommender system, potential manipulation by the platform, echo chamber effects, or lack of personalization for less active users due to the dominance of active users during collaborative learning. Therefore, there is an urgent need to develop a new paradigm to protect user interests and alleviate these issues. Recently, some researchers have introduced LLM agents to simulate user behaviors, these approaches primarily aim to optimize platform-side performance, leaving core issues in recommender systems unresolved. To address these limitations, we propose a new user-agent-platform paradigm, where agent serves as the protective shield between user and recommender system that enables indirect exposure. To this end, we first construct four recommendation datasets, denoted as InstructRec, along with user instructions for each record. To understand user’s intention, we design an Instruction-aware Agent (InstructAgent) capable of using tools to acquire knowledge from external environments. Moreover, we introduce an Individual Instruction-aware Agent (Instruct2Agent), which incorporates a dynamic memory mechanism to optimize from individual feedback. Results on four InstructRec datasets demonstrate that Instruct2Agent consistently achieves an average improvement of 16.6% over SOTA baselines across ranking metrics. Moreover, Instruct2Agent mitigates echo chamber effects and effectively alleviates the model bias in disadvantaged users (less-active), serving as a shield between user and recommender systems. Datasets and code are publicly available at the URL [^1] https://github.com/WujiangXu/iAgent.
Recommender System, LLM Agent and LLM
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paper
1. Introduction
@@
(a) Previous User-Platform Paradigm
(b) Ours User-Agent-Platform Paradigm
Figure 1. (a) Previous recommendation ecosystem primarily focused on designing sophisticated models to enhance the ranking performance so as to increase platform’s benefit. However, they overlooked the user’s proactive instructions and put users under the direct control of recommender systems. (b) In contrast, we build an individual instruction-aware agent for each user, which generates re-ranking results based on the user’s active instructions. The agent’s memory component is influenced solely by the individual user, providing an individualized personal service.
@@
2. Task Definitions and Comparisions
Table 1. Difference between previous recommendation models and our model.
SR N/A N/A
CRS Fixed Multiple Turns
RecAgent N/A N/A
Ours Flexible 0, 1, or Multiple Turns
(1) $\hat{i}=\arg\max_{i\in I}P(s_{T+1}=i\mid S_{u};\psi).$
(2) $\hat{i}=\arg\max_{i\in I}P(s_{T+1}=i\mid S_{u},H_{u};\psi).$
(3) $\hat{i}=\arg\max_{i\in I}P(s_{T+1}=i\mid S_{u},\Omega_{u},E;\psi_{u}).$
3. Methodology
In this part, we firstly introduce the naive solution InstructAgent based on InstructRec, which can learn the intention from the user instruction. Next, we introduce our Instruct2Agent equipped with individual dynamic memory. The workflow of models are shown in Fig. 2. All the prompt templates used in InstructAgent and Instruct2Agent and examples of responses are provided in Appendix C.
(a) InstructAgent
(b) Instruct2Agent
Figure 2. Workflow of our proposed agents. (a) InstructAgent explores the relative knowledge under the user’s instruction and provides the reranking results refined by the self-reflection mechanism. (b) Instruct2Agent designs the dynamic memory mechanism to improve the personalized ability of InstructAgent.
3.1. InstructAgent
@@
(5) $\mathcal{R^{*}}\leftarrow M_{r}(X_{IK}\,\|X_{EK}\,\|X_{SU}\,\|\,X_{Item}\,\|\,P_{tr})$
@@
Self-reflection Mechanism. Large language models output content in a generative manner, which can lead to hallucination problems (Huang et al., 2023b). To address this, we designed a self-reflection mechanism to verify the content of the re-ranked item list. Specifically, we compare the elements between the reranking list and the previous one. If no differences are found, the results are directly output. However, if discrepancies are detected, the self-reflection module invokes the reranker to regenerate the reranking list, adding a prompt $P_{sr}$ to ensure alignment with the original ranked list. The mathematical formulation remains the same as in Eq. 5, with the prompt $P_{tr}$ replaced by $P_{sr}$.
3.2. Instruct2Agent
Although our basic framework InstructAgent can explore knowledge based on the user’s instructions, it fails to effectively model the dynamic interests within the instructions and cannot learn from user feedback.. To address this, we design a profile generator to build user’s personal profile that learns from the user feedback and a dynamic extractor to extract dynamic interest and build dynamic profile according to the instruction. Unlike existing recommendation models, Instruct2Agent is uniquely optimized for individual users, remaining unaffected by the behaviors of other users.
@@ 3.2. Instruct2Agent
(7) $\mathcal{F}^{T}\leftarrow M_{ge}(\mathcal{F}^{T-1}\,\|X_{i}^{+*}\,\|X_{G}^{T}\,\|P_{pr2})$
@@
(8) $\mathcal{F}_{d}^{T}\,,X_{DU}\leftarrow M_{e}(\mathcal{F}^{T}\,\|X_{SU}\,\|X_{I}\,\|X_{IK}\,\|X_{EK}\,\|P_{e})$
@@
(9) $\mathcal{R^{*}}\leftarrow M_{r}(X_{IK}\,\|X_{EK}\,\|X_{SU}\,\|\,\mathcal{F}_{d}^{T}\,\|\,X_{DU}\,\|\,X_{Item}\,\|\,P_{tr}^{*})$
where $P_{tr}^{*}$ represents the prompt template for the reranker in Instruct2Agent. Besides, a self-reflection mechanism is also implemented to ensure consistent results, using the same inputs as the reranker, except for the prompt template.
4. Empirical Evaluation
In this section, we present extensive experiments to demonstrate the effectiveness of InstructAgent and Instruct2Agent, aiming to answer the following four research questions (RQs).
$\bullet$ RQ1: How does the performance of InstructAgent and Instruct2Agent compare to state-of-the-art baselines across various datasets?
@@ 4.1. Experiment Setup
Instruction Generator: Initially, we manually annotate several instruction-review pairs, providing few-shot examples for LLMs to facilitate in-context learning. These few-shot examples, along with reviews paired with a random persona from Persona Hub (Chan et al., 2024), are then fed into the LLM [^4] We use GPT-4o-mini for data generation. to generate instructions. To ensure that the few-shot examples remain dynamic, we create a list to store the instruction-review pairs and allow the LLM to decide whether a newly generated instruction should be included as an example. Examples of the annotated instruction-review pairs, generated instructions, and the data construction processes can be found in Appendix C.3.
Instruction Cleaner: To prevent data leakage from the reviews, we test if or not the LLM can recover the item from the generated instruction. More specifically, given the instruction, we employ the LLM to choose between the ground-truth item and a randomly selected negative item. The LLM generates a certainty score based on the instruction and the item’s textual information. Based on the result, we retain all of those instructions for which the LLM cannot infer the ground-truth item, and also keep an equal number of correctly inferred instructions that has low certainty scores. Statistical analysis of InstructRec dataset is in Table 2. For the filtered instructions and the retained instructions, we show some examples in Appendix C.3.2.
Figure 3. The overview of our InstructRec dataset construction.
Table 2. Statistics of the InstructRec dataset: $\left|\mathcal{U}\right|$, $\left|\mathcal{V}\right|$, and $\left|\mathcal{E}\right|$ represent the number of users, items, and interactions, respectively. #$\left|X_{I}\right|$ denotes the average token length of user instructions, while #$\left|S_{U}\right|$ represents the average token length of the user’s static memory.
@@
Evaluation Protocol. We randomly sample 9 negative items with one true item to make the candidate ranking list. Following the data split in sequential recommendation (Kang and McAuley, 2018), the most recent interaction is reserved for testing. The agent-based works, including ours, utilize all the interaction data except the most recent one to construct the agent’s memory. For evaluation metric, we adopt the typical top-$N$ metrics hit rate (HR@{1, 3}), normalized discounted cumulative gain (NDCG@{3}) (Järvelin and Kekäläinen, 2002) and Mean Reciprocal Rank (MRR) (Sarwar et al., 2001). In addition to conventional ranking metrics, we conduct additional experiments to ensure that our InstructAgent/Instruct2Agent can act as a shield between users and the recommendation system. Specifically, we design evaluation metrics such as the percentage of filtered Ads items (FR@1,3,5,10) and popularity-weighted ranking metrics (P-HR@3 and P-MRR) to validate the mitigation of the echo chamber effect (Ge et al., 2020; Xu et al., 2022). We use freqi to denote the frequency of item $i$ in the dataset. Formally, these metrics are defined as:
@@
Baselines. We compare our method with three classes of baselines: (1) Sequential recommendation methods, i.e., BERT4Rec (Sun et al., 2019), GRU4Rec (Hidasi et al., 2015) and SASRec (Kang and McAuley, 2018). (2) Instruction-aware methods, i.e., BM25 (Robertson et al., 2009), BGE-Rerank (Xiao et al., 2023) and EasyRec (Ren and Huang, 2024). (3) Recommendation agents, i.e., ToolRec (Zhao et al., 2024) and AgentCF (Zhang et al., 2024b). Detailed implementation and introduction of baselines are in Appendix B.
4.2. Performance Comparison
Main Results. (RQ1) Tables 3 and 4 present the experimental results across four datasets using different evaluation metrics. By incorporating instruction knowledge into the model, the instruction-aware baselines outperform traditional recommendation agent methods. Benefiting from the alignment with collaborative filtering and natural language information, EasyRec pretraining on several Amazon datasets achieves the second-best results, trailing only our InstructAgent. Our Instruct2Agent outperforms the second-best baseline, EasyRec, with the averagely 16.6% improvement. This improvement is partly attributed to the parser component, which learns instruction-aware knowledge, enabling the reranker to better understand the user’s intentions. Meanwhile, our proposed dynamic memory component leverages user feedback to construct a more accurate user profile and dynamically extract interests from historical data based on the instruction.
Table 3. Evaluation results (%) of the ranking metric ($\uparrow$) on the InstructRec. We highlight the methods with the first, second and third best performances.
@@ 4.2. Performance Comparison
InstructAgent 31.89 48.99 41.69 47.23 38.19 56.87 48.93 53.04
Instruct2Agent 35.11 53.51 45.64 50.28 46.43 65.77 57.67 60.43
Table 4. Evaluation results (%) of the ranking metric ($\uparrow$) on InstructRec.
@@
InstructAgent 23.56 47.01 36.98 42.19 37.40 56.33 48.28 52.42
Instruct2Agent 30.97 56.69 45.76 49.14 39.22 57.92 49.96 53.78
Echo Chamber Effect. (RQ2) We also report the experimental results evaluating the echo chamber effect in Table 5. Ads items are randomly inserted into the candidate ranking list from other domains to simulate advertising scenarios that users may have encountered. To mitigate position bias in LLMs (Liu et al., 2024a), Ads items are added randomly within the candidate list positions. Instruct2Agent accurately identifies users’ instructions and extracts knowledge about their underlying needs, thereby effectively removing undesired Ads. Benefitting from not being trained in a purely data-driven manner and constructing user profiles based on their feedback, our Instruct2Agent also recommends more diverse items to users (both active and less-active items), instead of focusing solely on popular items, and meanwhile improves the overall recommendation performance. Drawing from these experimental results, we conclude that our Instruct2Agent can mitigate the echo chamber effect and act as a protective shield for users. Due to the page limitation, we provide full experiment results in Appendix B.3.1.
Table 5. Evaluation of the echo chamber effects (%) ($\uparrow$) on InstructRec.
@@
InstructAgent 71.98 67.82 59.51 57.32 78.24 69.71 62.74 58.76
Instruct2Agent 77.15 70.15 64.70 60.87 87.69 84.20 64.48 60.20
Protect Less-Active Users. (RQ3) We define the top 20% of users as active, with the remaining 80% classified as less-active (Li et al., 2021; Xu et al., 2023a). Since our data is sampled and filtered using a 10-core process, most users exhibit rich behavioral patterns. Consequently, active users tend to show poorer performance compared to less-active users, largely due to the decline in LLM performance with longer texts (Liu et al., 2024b). As illustrated in Table 6, our Instruct2Agent enhances the performance for both active and less-active users. For less-active users, we construct individual profiles based on their feedback, ensuring that these profiles are not influenced by other users. The experimental results demonstrate that our dynamic memory mechanism offers personalized services tailored to each user individually. Detailed implementation and introduction of baselines are in Appendix B.3.2.
Table 6. The performance (%) of active and less-active users on InstructRec - Amazon book.
@@
InstructAgent 34.07 50.79 43.67 49.00 29.96 47.73 40.14 45.71
Instruct2Agent 37.92 55.75 47.84 52.11 33.27 51.74 43.81 48.67
Model Study. (RQ4) First, we analyze the impact of our self-reflection mechanism on the LLM’s hallucination rate. When implementing ToolRec (Zhao et al., 2024) and AgentCF (Zhang et al., 2024b), we applied the self-reflection mechanism to improve the accuracy of the reranking list. As shown in Fig. 4, the self-reflection mechanism reduces the hallucination rate by at least 20-fold. In this mechanism, we prompt the LLM to generate the reranking list based on the initial ranking list. However, Instruct2Agent exhibits the highest error rate, as the longer text sequence causes the LLM to lose some information from the original ranking list. Based on the experimental results, we can safely conclude that our self-reflection mechanism effectively alleviates LLM-induced hallucinations.
@@
Figure 4. The first row presents the hallucination rate with and without the self-reflection mechanism, while the second row illustrates the probability of changes in the ranking list after our reranker.
5. Related Work
5.1. Recommender System
Sequential recommendation methods (Hidasi et al., 2015; Sun et al., 2019; Kang and McAuley, 2018; Ning et al., 2024) primarily focus on developing temporal encoders to capture both short- and long-term user interests. The evolution of these encoders has progressed from GRU units (Hidasi et al., 2015), to more advanced architectures such as self-attention mechanisms (e.g., SASRec (Kang and McAuley, 2018)), bidirectional encoders (e.g., BERT4Rec (Sun et al., 2019) with masked item training), graph neural networks (He et al., 2020; Xu et al., 2023b, 2024c), and other Transformer-based models (Xu et al., 2024a). In the context of embracing large language models, generative recommenders (Geng et al., 2022; Mei and Zhang, 2023; Zhai et al., 2024) treat item indices as tokens and predict them in a generative manner. Meanwhile, LLMs (Li et al., 2023; Xu et al., 2024b) are utilized to play as a sequential embedding extractor to improve the recommendation performance. In our framework design, all recommendation models can be considered as components of the tools.

### v2

# iAgent: LLM Agent as a Shield between User and Recommender Systems

@@
Abstract
Traditional recommender systems usually take the user-platform paradigm, where users are directly exposed under the control of the platform’s recommendation algorithms. However, the defect of recommendation algorithms may put users in very vulnerable positions under this paradigm. First, many sophisticated models are often designed with commercial objectives in mind, focusing on the platform’s benefits, which may hinder their ability to protect and capture users’ true interests. Second, these models are typically optimized using data from all users, which may overlook individual user’s preferences. Due to these shortcomings, users may experience several disadvantages under the traditional user-platform direct exposure paradigm, such as lack of control over the recommender system, potential manipulation by the platform, echo chamber effects, or lack of personalization for less active users due to the dominance of active users during collaborative learning. Therefore, there is an urgent need to develop a new paradigm to protect user interests and alleviate these issues. Recently, some researchers have introduced LLM agents to simulate user behaviors, these approaches primarily aim to optimize platform-side performance, leaving core issues in recommender systems unresolved. To address these limitations, we propose a new user-agent-platform paradigm, where agent serves as the protective shield between user and recommender system that enables indirect exposure. To this end, we first construct four recommendation datasets, denoted as InstructRec, along with user instructions for each record. To understand user’s intention, we design an Instruction-aware Agent (iAgent) capable of using tools to acquire knowledge from external environments. Moreover, we introduce an Individual Instruction-aware Agent (i2Agent), which incorporates a dynamic memory mechanism to optimize from individual feedback. Results on four InstructRec datasets demonstrate that i2Agent consistently achieves an average improvement of 16.6% over SOTA baselines across ranking metrics. Moreover, i2Agent mitigates echo chamber effects and effectively alleviates the model bias in disadvantaged users (less-active), serving as a shield between user and recommender systems. Datasets and code are publicly available at the URL [^1] https://anonymous.4open.science/r/iAgent-8F65.
iAgent: LLM Agent as a Shield
between User and Recommender Systems
Wujiang Xu1, Yunxiao Shi2, Zujie Liang3, Xuying Ning4, Kai Mei1, Kun Wang5, Xi Zhu1, Min Xu2, Yongfeng Zhang1 1Rutgers University, 2University of Technology Sydney, 3Independent Researcher, 4University of Illinois Urbana-Champaign, 5Nanyang Technological University wujiang.xu@rutgers.edu, yongfeng.zhang@rutgers.edu
1 Introduction
@@
Figure 1: (a) Previous recommendation ecosystem primarily focused on designing sophisticated models to enhance the ranking performance so as to increase platform’s benefit. However, they overlooked the user’s proactive instructions and put users under the direct control of recommender systems. (b) In contrast, we build an individual instruction-aware agent for each user, which generates re-ranking results based on the user’s active instructions. The agent’s memory component is influenced solely by the individual user, providing an individualized personal service.
@@
2 Methodology
In this part, we firstly introduce the naive solution iAgent based on InstructRec, which can learn the intention from the user instruction. Next, we introduce our i2Agent equipped with individual dynamic memory. The workflow of models are shown in Fig. 2. All the prompt templates used in iAgent and i2Agent and examples of responses are provided in Appendix D.
(a) iAgent
(b) i2Agent
Figure 2: Workflow of our proposed agents. (a) iAgent explores the relative knowledge under the user’s instruction and provides the reranking results refined by the self-reflection mechanism. (b) i2Agent designs the dynamic memory mechanism to improve the personalized ability of iAgent.
2.1 iAgent
@@ 2 Methodology
$\mathcal{R^{*}}\leftarrow M_{r}(X_{IK}\,\|X_{EK}\,\|X_{SU}\,\|\,X_{Item}\,\|\,P_{tr})$ (2)
@@
Self-reflection Mechanism. Large language models output content in a generative manner, which can lead to hallucination problems (Huang et al., 2023a). To address this, we designed a self-reflection mechanism to verify the content of the re-ranked item list. Specifically, we compare the elements between the reranking list and the previous one. If no differences are found, the results are directly output. However, if discrepancies are detected, the self-reflection module invokes the reranker to regenerate the reranking list, adding a prompt $P_{sr}$ to ensure alignment with the original ranked list. The formulation remains the same as in Eq. 2, with the prompt $P_{tr}$ replaced by $P_{sr}$.
2.2 i2Agent
Although our basic framework iAgent can explore knowledge based on the user’s instructions, it fails to effectively model the dynamic interests within the instructions and cannot learn from user feedback.. To address this, we design a profile generator to build user’s personal profile that learns from the user feedback and a dynamic extractor to extract dynamic interest and build dynamic profile according to the instruction. Unlike existing recommendation models, i2Agent is uniquely optimized for individual users, remaining unaffected by the behaviors of other users.
@@
$\mathcal{F}^{T}\leftarrow M_{ge}(\mathcal{F}^{T-1}\,\|X_{i}^{+*}\,\|X_{G}^{T}\,\|P_{pr2})$ (4)
@@
$\mathcal{F}_{d}^{T}\,,X_{DU}\leftarrow M_{e}(\mathcal{F}^{T}\,\|X_{SU}\,\|X_{I}\,\|X_{IK}\,\|X_{EK}\,\|P_{e})$ (5)
@@
$\mathcal{R^{*}}\leftarrow M_{r}(X_{IK}\,\|X_{EK}\,\|X_{SU}\,\|\,\mathcal{F}_{d}^{T}\,\|\,X_{DU}\,\|\,X_{Item}\,\|\,P_{tr}^{*})$ (6)
where $P_{tr}^{*}$ represents the prompt template for the reranker in i2Agent. Besides, a self-reflection mechanism is also implemented to ensure consistent results, using the same inputs as the reranker, except for the prompt template.
3 Empirical Evaluation
In this section, we present extensive experiments to demonstrate the effectiveness of iAgent and i2Agent, aiming to answer the following four research questions (RQs).
$\bullet$ RQ1: How does the performance of iAgent and i2Agent compare to state-of-the-art baselines across various datasets?
@@ 3.1 Experiment Setup
Instruction Generator: Initially, we manually annotate several instruction-review pairs, providing few-shot examples for LLMs to facilitate in-context learning. These few-shot examples, along with reviews paired with a random persona from Persona Hub (Chan et al., 2024), are then fed into the LLM [^3] We use GPT-4o-mini for data generation. to generate instructions. To ensure that the few-shot examples remain dynamic, we create a list to store the instruction-review pairs and allow the LLM to decide whether a newly generated instruction should be included as an example. Examples of the annotated instruction-review pairs, generated instructions, and the data construction processes can be found in Appendix D.3.
Instruction Cleaner: To prevent data leakage from the reviews, we test if or not the LLM can recover the item from the generated instruction. More specifically, given the instruction, we employ the LLM to choose between the ground-truth item and a randomly selected negative item. The LLM generates a certainty score based on the instruction and the item’s textual information. Based on the result, we retain all of those instructions for which the LLM cannot infer the ground-truth item, and also keep an equal number of correctly inferred instructions that has low certainty scores. Statistical analysis of InstructRec dataset is in Table 1. For the filtered instructions and the retained instructions, we show some examples in Appendix D.3.2.
Figure 3: The overview of our InstructRec dataset construction.
Table 1: Statistics of the InstructRec dataset: $\left|\mathcal{U}\right|$, $\left|\mathcal{V}\right|$, and $\left|\mathcal{E}\right|$ represent the number of users, items, and interactions, respectively. #$\left|X_{I}\right|$ denotes the average token length of user instructions, while #$\left|S_{U}\right|$ represents the average token length of the user’s static memory.
@@
Evaluation Protocol. We randomly sample 9 negative items with one true item to make the candidate ranking list. Following the data split in sequential recommendation (Kang and McAuley, 2018), the most recent interaction is reserved for testing. The agent-based works, including ours, utilize all the interaction data except the most recent one to construct the agent’s memory. For evaluation metric, we adopt the typical top-$N$ metrics hit rate (HR@{1, 3}), normalized discounted cumulative gain (NDCG@{3}) (Järvelin and Kekäläinen, 2002) and Mean Reciprocal Rank (MRR) (Sarwar et al., 2001). In addition to conventional ranking metrics, we conduct additional experiments to ensure that our iAgent/i2Agent can act as a shield between users and the recommendation system. Specifically, we design evaluation metrics such as the percentage of filtered Ads items (FR@1,3,5,10) and popularity-weighted ranking metrics (P-HR@3 and P-MRR) to validate the mitigation of the echo chamber effect (Ge et al., 2020; Xu et al., 2022). We use freqi to denote the frequency of item $i$ in the dataset. Formally, these metrics are defined as:
[…]
@@
Baselines. We compare our method with three classes of baselines: (1) Sequential recommendation methods, i.e., BERT4Rec (Sun et al., 2019), GRU4Rec (Hidasi et al., 2015) and SASRec (Kang and McAuley, 2018). (2) Instruction-aware methods, i.e., BM25 (Robertson et al., 2009), BGE-Rerank (Xiao et al., 2023) and EasyRec (Ren and Huang, 2024). (3) Recommendation agents, i.e., ToolRec (Zhao et al., 2024) and AgentCF (Zhang et al., 2024b). Detailed implementation and introduction of baselines are in Appendix C.
3.2 Performance Comparison
Main Results. (RQ1) Tables 2 and 3 present the experimental results across four datasets using different evaluation metrics. By incorporating instruction knowledge into the model, the instruction-aware baselines outperform traditional recommendation agent methods. Benefiting from the alignment with collaborative filtering and natural language information, EasyRec pretraining on several Amazon datasets achieves the second-best results, trailing only our iAgent. Our i2Agent outperforms the second-best baseline, EasyRec, with the averagely 16.6% improvement. This improvement is partly attributed to the parser component, which learns instruction-aware knowledge, enabling the reranker to better understand the user’s intentions. Meanwhile, our proposed dynamic memory component leverages user feedback to construct a more accurate user profile and dynamically extract interests from historical data based on the instruction.
Table 2: Evaluation results (%) of the ranking metric ($\uparrow$) on the InstructRec. We highlight the methods with the first, second and third best performances.
@@ 3.2 Performance Comparison
iAgent 31.89 48.99 41.69 47.23 38.19 56.87 48.93 53.04
i2Agent 35.11 53.51 45.64 50.28 46.43 65.77 57.67 60.43
Table 3: Evaluation results (%) of the ranking metric ($\uparrow$) on InstructRec.
@@
iAgent 23.56 47.01 36.98 42.19 37.40 56.33 48.28 52.42
i2Agent 30.97 56.69 45.76 49.14 39.22 57.92 49.96 53.78
Echo Chamber Effect. (RQ2) We also report the experimental results evaluating the echo chamber effect in Table 4. Ads items are randomly inserted into the candidate ranking list from other domains to simulate advertising scenarios that users may have encountered. To mitigate position bias in LLMs (Liu et al., 2024a), Ads items are added randomly within the candidate list positions. i2Agent accurately identifies users’ instructions and extracts knowledge about their underlying needs, thereby effectively removing undesired Ads. Benefitting from not being trained in a purely data-driven manner and constructing user profiles based on their feedback, our i2Agent also recommends more diverse items to users (both active and less-active items), instead of focusing solely on popular items, and meanwhile improves the overall recommendation performance. Drawing from these experimental results, we conclude that our i2Agent can mitigate the echo chamber effect and act as a protective shield for users. Due to the page limitation, we provide full experiment results in Appendix C.3.1.
Table 4: Evaluation of the echo chamber effects (%) ($\uparrow$) on InstructRec.
@@
iAgent 71.98 67.82 59.51 57.32 78.24 69.71 62.74 58.76
i2Agent 77.15 70.15 64.70 60.87 87.69 84.20 64.48 60.20
Protect Less-Active Users. (RQ3) We define the top 20% of users as active, with the remaining 80% classified as less-active (Li et al., 2021; Xu et al., 2023). Since our data is sampled and filtered using a 10-core process, most users exhibit rich behavioral patterns. Consequently, active users tend to show poorer performance compared to less-active users, largely due to the decline in LLM performance with longer texts (Liu et al., 2024b). As illustrated in Table 5, our i2Agent enhances the performance for both active and less-active users. For less-active users, we construct individual profiles based on their feedback, ensuring that these profiles are not influenced by other users. The experimental results demonstrate that our dynamic memory mechanism offers personalized services tailored to each user individually. Detailed implementation and introduction of baselines are in Appendix C.3.2.
Table 5: The performance (%) of active and less-active users on InstructRec - Amazon book.
@@
iAgent 34.07 50.79 43.67 49.00 29.96 47.73 40.14 45.71
i2Agent 37.92 55.75 47.84 52.11 33.27 51.74 43.81 48.67
Model Study. (RQ4) First, we analyze the impact of our self-reflection mechanism on the LLM’s hallucination rate. When implementing ToolRec (Zhao et al., 2024) and AgentCF (Zhang et al., 2024b), we applied the self-reflection mechanism to improve the accuracy of the reranking list. As shown in Fig. 4, the self-reflection mechanism reduces the hallucination rate by at least 20-fold. In this mechanism, we prompt the LLM to generate the reranking list based on the initial ranking list. However, i2Agent exhibits the highest error rate, as the longer text sequence causes the LLM to lose some information from the original ranking list. Based on the experimental results, we can safely conclude that our self-reflection mechanism effectively alleviates LLM-induced hallucinations.
@@
Figure 4: The first row presents the hallucination rate with and without the self-reflection mechanism, while the second row illustrates the probability of changes in the ranking list after our reranker.
4 Related Work
4.1 Recommender System
Sequential recommendation models (Hidasi et al., 2015) focus on developing temporal encoders to capture both short- and long-term user interests, exemplified by SASRec (Kang and McAuley, 2018)’s attention mechanism and BERT4Rec (Sun et al., 2019)’s bidirectional encoder. Recent developments have integrated large language models, with some approaches treating item indices as tokens for generative recommendations (Geng et al., 2022), while others utilize LLMs (Li et al., 2023; Xu et al., 2024) as sequential embedding extractors to enhance recommendation performance. The rise of LLMs has also transformed conversational recommendation systems (CRS)(Sun and Zhang, 2018; Zhang et al., 2018), improving dialogue understanding and flexibility(Friedman et al., 2023; Feng et al., 2023) compared to conventional approaches that were limited in dialogue format and turn count.
4.2 Personal Language-based Agent
With the advancement of large language models (Achiam et al., 2023), research has evolved from early persona-based dialogue agents (Zhang, 2018; Park et al., 2023) to more sophisticated domain-specific agents (Gur et al., 2023; Deng et al., 2024; Xie et al., 2024) incorporating tool learning and memory mechanisms. In the recommendation domain, recent work has developed recommendation agents (Zhao et al., 2024; Wang et al., 2023; Zhang et al., 2024a) that leverage historical interaction information as user memory and utilize LLMs for ranking. Notably, newer approaches focus on user-side operations (Wang et al., 2024; Huang et al., 2023b), generating re-ranking results based on user instructions and individual memory. We provide a more detailed related work in Appendix B.
5 Conclusion
In this work, we first design a straightforward instruction-aware agent (iAgent) to analyze user instructions and integrate relevant and comprehensive knowledge. Moreover, to enhance the agent’s personalized abilities, we propose individual instruction-aware agent (i2Agent), which incorporates a dynamic memory mechanism to learn from user’s personal feedback and extracts the dynamic interests. In addition to these technical contributions, our work also presents unique and complementary avenues for future research. We discuss potential future directions and open challenges in Appendix E.
6 Limitation
While our work demonstrates promising results, there are several limitations to note. First, our current implementation primarily focuses on English instructions, and the effectiveness of the model across different languages remains to be explored. recommendation scenarios. Additionally, while our evaluation metrics show improvements in recommendation quality, they may not fully capture the nuanced aspects of user satisfaction and long-term engagement. These limitations suggest potential directions for future research in developing more efficient and comprehensive instruction-aware recommendation systems.
APPENDIX
Appendix A Task Definitions and Comparisions
$\hat{i}=\arg\max_{i\in I}P(s_{T+1}=i\mid S_{u};\psi).$ (9)
$\hat{i}=\arg\max_{i\in I}P(s_{T+1}=i\mid S_{u},H_{u};\psi).$ (10)
$\hat{i}=\arg\max_{i\in I}P(s_{T+1}=i\mid S_{u},\Omega_{u},E;\psi_{u}).$ (11)
Table 6: Difference between previous recommendation models and our model.
SR N/A N/A
CRS Fixed Multiple Turns
RecAgent N/A N/A
Ours Flexible 0, 1, or Multiple Turns
Appendix B Detailed Related Work
B.1 Recommender System
Sequential recommendation models (Hidasi et al., 2015; Sun et al., 2019; Kang and McAuley, 2018) primarily focus on developing temporal encoders to capture both short- and long-term user interests. For instance, SASRec (Kang and McAuley, 2018) leverages an attention mechanism to capture long-term semantics, while BERT4Rec (Sun et al., 2019) uses a bidirectional encoder with a masked item training objective. In the context of embracing large language models, generative recommenders (Geng et al., 2022; Zhai et al., 2024) treat item indices as tokens and predict them in a generative manner. Meanwhile, LLMs (Li et al., 2023; Xu et al., 2024) are utilized to play as a sequential embedding extractor to improve the recommendation performance. In our framework design, all recommendation models can be considered as components of the tools.